What length of resistance of diameter 0.6mm and resistivity 1.1×10^-6 would u cut from a reel in order to make a 4.4 resistor


1
Expert's answer
2021-03-09T15:27:16-0500
"R=\dfrac{\rho L}{A}=\dfrac{4\rho L}{\pi d^2},""L=\dfrac{R\pi d^2}{4\rho},""L=\dfrac{4.4\ \Omega\cdot\pi\cdot(0.6\cdot10^{-3}\ m)^2}{4\cdot1.1\cdot10^{-6}\ \Omega\cdot m}=1.13\ m."